# Prominence-Aware Artifact Detection and Dataset for Image Super-Resolution

### v1

@@ 4.4 Training
We train our MLP prominence regressor using Adam on a training subset of 374 artifact examples from our dataset described in Section 3. The MLP predicts a prominence value for each pixel of the input image. We compute the mean predicted prominence inside and outside the binary artifact mask from the dataset. The training loss consists of two $L_{2}$ components:
$\mathcal{L}=L_{2}(\mathrm{MeanInside},\mathrm{GT\ Prominence})+L_{2}(\mathrm{MeanOutside},0).$ (1)
The model is trained to predict the ground-truth prominence value inside the binary mask, and 0 (no artifact) outside it. Thanks to small model size, the training converges quickly, usually in around 10–30 epochs. One training epoch takes about 13 seconds on an Nvidia RTX 3090 GPU.
@@ 5 Experiments
Our evaluation comprises three components. Section 5.1 provides an overview of our approach to objectively evaluating artifact-detection methods using our prominence dataset, then Section 5.2 compares our method to existing work with objective scores. Next, Section 5.3 evaluates methods on the primary downstream task: finding artifacts prominent to human viewers. Finally, Section 5.4 evaluates methods on the secondary downstream task: reducing SR proneness to artifacts.
@@ 5.1 Objective evaluation methodology
Our evaluation follows standard binary classification methodology—binarizing detection outputs with thresholds (selected for each method in evaluation) and computing TP, FP, FN—but modifies the precision and recall to weight detections by their prominence scores (Figure 7). Following Rachakonda & Bhatnagar (2021); Harju & Mesaros (2023), we treat artifact labels as graded values rather than binary. In our setting, each label is a prominence score rather than a probability of class membership. A low prominence score indicates that the masked area contains no artifact, so a positive detection in such cases should be penalized. We implement this using a margin $\kappa=0.3$, yielding:
$Prec^{pr}=\frac{\sum_{i\in\text{images}}TP_{i}\ast(\text{p}_{i}-\kappa)}{\sum_{i\in\text{images}}(TP_{i}+FP_{i})};\quad Rec^{pr}=\frac{\sum_{i\in\text{images}}TP_{i}\ast(\text{p}_{i}-\kappa)}{\sum_{i\in\text{images}}(TP_{i}+FN_{i})}.$ (2)
We set $\kappa$ = 0.3 to reward metrics that detect even less prominent artifacts: if more than 30% of viewers reported seeing the artifact, we already already consider it prominent enough to detect. This choice was guided by dataset statistics and observed viewer consistency.
From $Prec^{pr}$ and $Rec^{pr}$, we compute F1-score and PR-AUC. Note that unlike binary classification, scores computed this way have a different range from $[0,1]$ (see Appendix F). However, they remain suitable for ranking methods relatively to each other.
@@ 5.2 Objective prominence-metric evaluation
We evaluate our method using both our proposed methodology, and the methodology from DeSRA (Xie et al., 2023a), to avoid bias towards our own dataset and scoring. Following DeSRA, we selected binarization thresholds for all methods by maximizing the Precision × Recall product on the prominent subset of the DeSRA dataset. The thresholds are shown in all tables as t=0.xx.
Table 2 compares our method with other approaches on the basis of our prominence datasets, under different choices of reference input as described in Section 4.3. These include original high-resolution frames, the lightweight SR models SPAN (Hu et al., 2025) and RLFN (Kong et al., 2022), and, for DeSRA where HR frames are unavailable, the authors’ MSE-SR model trained with an MSE loss. Since the MSE-SR weights were not released, this reference cannot be used in other experiments.
Our method delivers the most consistent performance across all experiment settings, as reflected by its best average ranking in PR-AUC. These results also confirm that using pseudo-GT remains practical for artifact detection. We adopt RLFN pseudo-GT for all other experiments in the paper.
@@
We also evaluated our method on a learning-based compression task, as described in Appendix G. You can find $Prec^{pr}$ and $Rec^{pr}$ values for all methods in Section H.8.
@@ 5.3 Evaluating robustness to SR artifacts
SR Type Masks Found Mean Prominence Conf. Masks Found
DRCT (Hsu et al., 2024a) Transformer 43 11.85% 1
@@
Table 5 shows the results grouped by SR model. DRCT (Hsu et al., 2024a) and HAT-L (Chen et al., 2023), both Transformer based, show excellent results with nearly zero prominent artifacts detected. Next are three diffusion-based methods: SinSR (Wang et al., 2024b), ResShift (Yue et al., 2023), and StableSR (Wang et al., 2024a). Then, Table 5 shows the results grouped by artifact-detection method. Ours shares first place with DISTS in number of confident masks found while beating it in mean artifact prominence. DeSRA ranked third in this evaluation.
Also note that LDL with a 0.005 threshold finds regions with highly visible artifacts, but it trails far behind other methods in total number of confident masks. To account for both of these scores, we multiplied them analogously to the precision × recall product of Xie et al. (2023a); the results are in the last column. We tested LDL at two lower thresholds, which increased the total masks found, but they mainly captured non-prominent artifacts, yielding worse combined score.
Appendix E contains extra crowd-sourced evaluation results on other SR datasets.
Target SR LDL Liang et al. (2022) RealESRGAN Wang et al. (2021b) SwinIR Liang et al. (2021) Avg.
[…]
LDL (t=0.005) 17.26 90.43 6.91 18.32 88.94 11.56 4.22 77.94 21.08 4.9
LPIPS (t=0.25) 25.18 15.43 19.68 33.66 8.54 34.17 11.21 2.94 53.43 2.2
ERQA (t=0.55) 0.67 100.00 0.00 1.22 99.50 0.00 0.19 98.04 0.00 6.0
DeSRA 29.18 54.26 25.00 33.66 34.17 56.78 8.98 32.84 28.43 2.9
Ours (t=0.3) 34.71 20.74 45.21 38.01 14.57 55.78 11.86 6.37 57.35 1.6
@@ 5.4 Fine-tuning SR models to reduce artifacts
We follow the methodology from Xie et al. (2023a) and fine-tune SR models to reduce artifacts using artifact-detection methods. We fine-tune each SR with each method and compare the results.
Fine-tuning optimizes the pixel-wise MSE loss between the SR model’s output and an artificial GT image, constructed by replacing the artifact regions on the SR model’s output with output from RLFN. The artifact regions are the binarized output masks of a given artifact-detection method.
@@ 6 Conclusion
In this work, we address the challenge of visual artifacts in single-image super-resolution, a problem that plagues even the newest, most capable models such as Yu et al. (2024). We argued that artifacts should be characterized by their prominence to human observers rather than by binary masks, as the perceptual impact of artifacts varies significantly. We validated this perspective by showing that most viewers fail to notice almost half of the artifacts in the DeSRA dataset, which was annotated in-lab.
Our primary contribution is a new dataset of 1302 artifact examples with prominence annotations from 11 contemporary SISR methods. We also provide prominence annotations for all 593 artifacts from the DeSRA dataset.
Building on this, we developed an MLP-based method for SR-artifact detection that outputs prominence heatmaps, moving beyond simple binary masks. We conducted broad evaluations showing that compared with other methods, ours can more accurately evaluate the artifact severity of SR models. We showed the practical utility of our method in detecting prominent artifacts and fine-tuning SR models for artifact suppression.
@@
The implications of our work extend beyond SISR. The concept of artifact prominence is likely applicable to other image processing and restoration tasks, as suggested by our preliminary findings on a dataset of JPEG AI artifacts. Furthermore, prominence-aware metrics could guide future SR research to focus on structured regions where artifacts are most visible.
@@
We acknowledge the limitations of our study. The artifact masks in our dataset are approximate, as delineating exact artifact boundaries is ambiguous even for human annotators. Instead, we seed masks using existing methods, introducing some inaccuracy. Additionally, our proposed pseudo-GT approach relying on a lightweight SR model can lead to false positive detections when it fails to reconstruct fine textures that a more powerful model resolves correctly. Finally, the lack of existing prominence datasets limits our ability to faithfully validate our method.
There are several promising directions for future research. A natural extension is towards video super-resolution: while our work applies on a per-frame basis, we do not address temporal artifacts such as flickering. To mitigate the failure mode of a single lightweight SR, using an ensemble of models for pseudo-GT could be explored. Finally, higher-capacity models such as Yu et al. (2024) start to shift from simple texture distortions to more semantic artifacts like object replacement, inviting further investigation.
Our code and datasets, with prominence annotations, are available at tinyurl.com/2u9zxtyh.
Acknowledgments
This work was supported by the The Ministry of Economic Development of the Russian Federation in accordance with the subsidy agreement (agreement identifier  [^00] C313925P4H0002; grant No 139-15-2025-012).
The research was carried out using the MSU-270 supercomputer of Lomonosov Moscow State University.
Reproducibility statement
We have made our code, our artifact prominence dataset, and the prominence annotations for the DeSRA dataset publicly available, linked in Sections 1 and 6. The paper provides a detailed description of the dataset collection and crowdsourced assessment methodology in Section 3, the architecture and features of the proposed artifact detection method in Section 4, and the evaluation methodology in Section 5. Appendix B further describes the metrics used in the SR fine-tuning experiment.
Appendix A Crowdsourced annotation dispersion analysis
@@
Appendix B SR fine-tuning scoring and dilation details
Figure 9: Examples of SR artifacts removed after fine-tuning with the proposed metric, along with corresponding detected-artifact masks. The DeSRA dataset images are under the ImageNet license.
Figure 10: Schematic view of our SR fine-tuning pipeline.
@@ Appendix B SR fine-tuning scoring and dilation details
Section 5.4 overviews our fine-tuning process to reduce artifacts of existing SR models. Figure 9 shows fine-tuning examples. Here, we provide more details on the pipeline and on the scoring process.
Figure 10 shows our fine-tuning pipeline. The input low-resolution image is upscaled with the target SR and RLFN models, and the results are passed to the artifact metric. Then, the regions on the target-SR output where the artifacts were detected are replaced with regions from the RLFN output. The resulting artificial GT image is then used as a target to fine-tune the target-SR model.
@@
  • $\text{Rem}_{img}$ and $\text{Add}_{img}$, image-wise removal and addition rates. The ratio of images where $\left(\left|A\cap B\right|=0\right)\wedge\left(\left|B\right|\neq 0\right)$ determines whether the artifact was removed, and the ratio where $A\cup B>B$ determines whether a new artifact was introduced.
@@
  • $\text{Rem}_{pix}$. Pixel-wise removal rate; represents average number of pixels that were previously classified as artifacts and that go undetected after fine-tuning: $\left|\overline{A}\cap B\right|/\left|B\right|$.
@@
Appendix C Subjective quality of images with artifacts
@@
Appendix D Artifact examples and failure cases
@@
Figure 12 shows examples of prominent artifacts detected by our proposed method across various SR models (Yu et al., 2024; Wang et al., 2021b; 2024a; Liang et al., 2021; Wang et al., 2021a). Each example is annotated with the binary artifact mask and subjective prominence.
Figure 13 shows examples of false detections by our proposed method across SR models (Chen et al., 2023; Hsu et al., 2024b; Yue et al., 2023; Wang et al., 2024b; Cai et al., 2019). We observed the following failure cases:
@@
Appendix E Subjective evaluation on additional SR datasets

### v2

@@ 4.4 Training
We train our fusion module using Adam on a training subset of 374 artifact examples from our dataset described in Section 3. The model predicts a prominence value for each pixel of the input image. We compute the mean predicted prominence inside and outside the binary artifact mask from the dataset. The training loss consists of two $L_{2}$ components:
$\begin{split}\mathcal{L}&=L_{2}(\mathrm{MeanInside},\mathrm{GT\ Prominence})\\ &+L_{2}(\mathrm{MeanOutside},0).\end{split}$ (1)
The model is trained to predict the ground-truth prominence value inside the binary mask, and 0 (no artifact) outside it. Thanks to small model size, the training converges quickly, usually in around 10–15 epochs. One training epoch takes about 13 seconds on an Nvidia RTX 3090 GPU.
@@ 5 Experiments
Our evaluation comprises multiple steps. Section 5.1 provides an overview of our approach to objectively evaluating artifact-detection methods using our prominence dataset, then Section 5.2 compares our method to existing work with objective scores. Section 5.3 validates our method’s generalization across SR architectures. Next, Section 5.4 evaluates methods on the primary downstream task: finding artifacts prominent to human viewers. Finally, Section 5.5 evaluates methods on the secondary downstream task: reducing SR proneness to artifacts.
@@ 5.1 Objective evaluation methodology
Our evaluation follows the standard binary classification methodology—binarizing detection outputs with thresholds (selected for each method in evaluation) and computing TP, FP, FN—but modifies the precision and recall to weight detections by their prominence scores (Figure 7). Following (Rachakonda and Bhatnagar, 2021; Harju and Mesaros, 2023), we treat artifact labels as graded values rather than binary. In our setting, each label is a prominence score rather than a probability of class membership. A low prominence score indicates that the masked area contains no artifact, so a positive detection in such cases should be penalized. We implement this using a margin $\kappa=0.3$, yielding:
$\begin{split}Prec^{pr}&=\frac{\sum_{i\in\text{images}}TP_{i}\ast(\text{p}_{i}-\kappa)}{\sum_{i\in\text{images}}(TP_{i}+FP_{i})};\\ Rec^{pr}&=\frac{\sum_{i\in\text{images}}TP_{i}\ast(\text{p}_{i}-\kappa)}{\sum_{i\in\text{images}}(TP_{i}+FN_{i})}.\end{split}$ (2)
We set $\kappa=0.3$ to reward detection of less-prominent artifacts: if more than 30% of viewers reported the artifact, we consider it prominent enough to detect. This choice was guided by dataset statistics and observed viewer consistency.
From $Prec^{pr}$ and $Rec^{pr}$, we compute F1-score and PR-AUC. Unlike standard binary classification, these scores may fall outside $[0,1]$ (see Appendix G). However, they remain suitable for ranking methods relative to each other.
@@ 5.2 Objective prominence-metric evaluation
We evaluate our method using both our proposed methodology, and the methodology from DeSRA (Xie et al., 2023), to avoid bias towards our own dataset and scoring. Following DeSRA, we selected binarization thresholds for all methods by maximizing the Precision × Recall product on the prominent subset of the DeSRA dataset. The thresholds are shown in all tables as t=0.xx.
Table 2 compares our method with other approaches on the basis of our prominence datasets, under different choices of reference input as described in Section 4.3. These include original high-resolution frames, the lightweight SR models SPAN (Wan et al., 2024) and RLFN (Kong et al., 2022), and, for DeSRA where HR frames are unavailable, the authors’ MSE-SR model trained with an MSE loss. Since the MSE-SR weights were not released, this reference cannot be used in other experiments.
Our method delivers the most consistent performance across all experiment settings, as reflected by its best average ranking in PR-AUC. These results also confirm that pseudo-GT remains practical for artifact detection. We adopt RLFN pseudo-GT for all other experiments in the paper.
@@
We also evaluated our method on a learning-based compression task, as described in Appendix H. You can find $Prec^{pr}$ and $Rec^{pr}$ values for all methods in Section I.8.
Table 4: Cross-validation results on the prominent subset.
Val. Family CNN Transformer Diffusion
Method IoU PR-AUC IoU PR-AUC IoU PR-AUC
LDL (t=0.005) 0.100 0.258 0.027 0.154 0.133 0.256
SSIM (t=0.55) 0.218 0.416 0.236 0.337 0.222 0.530
LPIPS (t=0.25) 0.138 0.386 0.054 0.334 0.150 0.496
ERQA (t=0.55) 0.111 0.131 0.081 0.112 0.114 0.167
PaQ-2-PiQ (t=65) 0.031 0.027 0.023 0.037 0.023 0.031
TOPIQ (t=0.5) 0.173 0.106 0.143 0.088 0.052 0.100
DISTS (t=0.25) 0.187 0.292 0.165 0.299 0.291 0.416
bd_jup (t=0.1) 0.129 0.318 0.168 0.242 0.177 0.344
ssm_jup (t=0.2) 0.158 0.278 0.103 0.174 0.280 0.309
DeSRA (t=0.3) 0.158 0.372 0.063 0.224 0.170 0.424
Ours T+D (t=0.25) 0.239 0.470
Ours C+D (t=0) 0.231 0.419
Ours C+T (t=0.5) 0.331 0.482
5.3 Evaluating generalization across SR architectures
To evaluate our method’s ability to generalize across different SR models, we conduct a cross-validation experiment where we train on artifact examples from two SR families and validate on a held-out third family. We partition our artifact dataset by the SR family (CNN, transformer, or diffusion), train three separate models each excluding one family, and evaluate them on the excluded family’s artifacts.
Table 4 shows the results, including existing methods for comparison. Our models show the most consistent performance demonstrating that the learned artifact representations transfer across SR architectures. This suggests that artifacts from different SR families share common perceptual characteristics that our method successfully captures.
@@ 5.4 Evaluating robustness to SR artifacts
SR Family Masks Found Mean Prominence Conf. Masks Found
DRCT (Hsu et al., 2024) Transformer 43 11.85% 1
@@
Target SR LDL (Liang et al., 2022) RealESRGAN (Wang et al., 2021b) SwinIR (Liang et al., 2021) Avg.
[…]
LDL (t=0.005) 17.26 90.43 6.91 18.32 88.94 11.56 4.22 77.94 21.08 4.9
LPIPS (t=0.25) 25.18 15.43 19.68 33.66 8.54 34.17 11.21 2.94 53.43 2.2
ERQA (t=0.55) 0.67 100.00 0.00 1.22 99.50 0.00 0.19 98.04 0.00 6.0
DeSRA 29.18 54.26 25.00 33.66 34.17 56.78 8.98 32.84 28.43 2.9
Ours (t=0.3) 34.71 20.74 45.21 38.01 14.57 55.78 11.86 6.37 57.35 1.6
@@
Table 5 shows the results grouped by SR model. DRCT (Hsu et al., 2024) and HAT-L (Chen et al., 2023), both Transformer based, show excellent results with nearly zero prominent artifacts detected. Next are three diffusion-based methods: SinSR (Wang et al., 2024b), ResShift (Yue et al., 2023), and StableSR (Wang et al., 2024a). Then, Table 7 shows the results grouped by artifact-detection method. Ours shares first place with DISTS in number of confident masks found while beating it in mean artifact prominence. DeSRA ranked third in this evaluation. Appendix F provides extra crowd-sourced evaluation on other SR datasets.
Note that LDL with a 0.005 threshold finds highly visible artifacts, but it trails far behind other methods in total number of confident masks. To account for both of these scores, we multiplied them analogously to the precision × recall product of Xie et al. (2023); the results are in the last column. We tested LDL at two lower thresholds, which increased the total masks found, but they mainly captured non-prominent artifacts, yielding worse combined score.
@@ 5.5 Fine-tuning SR models to reduce artifacts
Following the methodology from Xie et al. (2023), we fine-tune SR models to reduce artifacts using artifact-detection methods, and compare all SR-detector combinations. Fine-tuning optimizes the pixel-wise MSE loss between the SR model’s output and an artificial GT image, constructed by replacing the artifact regions on the SR model’s output with output from RLFN. The artifact regions are the binarized output masks of a given artifact-detection method.
@@
Figure 8 demonstrates the resulting SR artifact reduction.
Figure 8: Examples of SR artifacts removed after fine-tuning with the proposed metric. Zoom in to see in full resolution.
@@ 6 Conclusion
We addressed the challenge of visual artifacts in single-image super-resolution by introducing the concept of prominence—characterizing artifacts by their perceptual impact rather than binary masks. Our crowdsourced study revealed that nearly half the artifacts in the existing DeSRA dataset go unnoticed by most viewers, validating this perspective. We contribute a new dataset of 1302 artifact examples with prominence annotations from 11 contemporary SISR methods, along with prominence labels for all 593 DeSRA artifacts. Building on this data, we developed a lightweight method that outputs prominence heatmaps and demonstrated its utility for evaluating SR models and fine-tuning them for artifact suppression.
@@
The implications of our work extend beyond SISR. Our evaluation on JPEG AI artifacts in Appendix H suggests that prominence modeling applies to other restoration tasks. Further, prominence-aware metrics could guide SR research toward structured regions where artifacts are most visible.
@@
We acknowledge limitations: the artifact masks in our dataset are approximate, as delineating exact artifact boundaries is ambiguous even for human annotators. Additionally, our pseudo-GT approach can produce false positives when the lightweight SR model fails to reconstruct fine textures.
Future work could extend prominence modeling to video super-resolution, addressing temporal artifacts such as flickering. Additionally, new higher-capacity models like SUPIR start to produce semantic artifacts (e.g., object replacement) rather than simple texture distortions—a shift that warrants further investigation.
Impact Statement
This paper presents work whose goal is to advance the field of image and video super-resolution. We aim to improve the robustness of SR methods to artifacts, making the output images appear more realistic with no disturbing details. Meanwhile, we make no attempt to prevent SR methods from generating natural-looking but incorrect details. This may be of critical importance to certain applications like face or license plate recognition, but falls outside the scope of our work.
Appendix A Input feature details
This section provides implementation details for the input features described in Section 4.1.
A.1 Structure Similarity Map (ssm_jup)
Our ssm_jup feature adapts the small-color-artifact detector from Tsereh et al. (2024), which is itself based on LDL (Liang et al., 2022). Given a reference image $I_{\text{ref}}$ (the original HR or pseudo-GT), the SR output $I_{\text{SR}}$, and a bicubic-upscaled baseline $I_{\text{bic}}$, we compute scaled residual-variance maps for both SR and bicubic outputs.
First, we compute the absolute residual summed across channels:
$R^{C}_{x}(i,j)=\sum_{c\in C}\bigl|I_{x}^{c}(i,j)-I_{\text{ref}}^{c}(i,j)\bigr|,\quad x\in\{\text{SR},\text{bic}\}.$ (3)
We then compute local variance within an $n\times n$ window and scale by a global factor:
$M^{C}_{x}(i,j)$ $=var\bigl(R^{C}_{x}\left(i-\frac{n-1}{2}:i+\frac{n-1}{2},j-\frac{n-1}{2}:j+\frac{n-1}{2}\right)\bigr),$ (4)
$S^{C}_{x}(i,j)$ $=var\bigl(R^{C}_{x}\bigr)^{1/5}\cdot M^{C}_{x}(i,j),$ (5)
where $n=33$.
The key modification from Tsereh et al. (2024) is in the choice of color channels $C$: the original method computes separate maps on chrominance channels (UV from YUV and ab from Lab color spaces) and intersects the thresholded results to detect color artifacts. We instead operate on all three RGB channels ($C=\{R,G,B\}$), which enables detection of luminance-correlated texture distortions that are common in SR artifacts but would be missed by chrominance-only analysis.
The final feature is the smoothed difference between the SR and bicubic maps:
$\text{ssm\_jup}=G_{\sigma}*S_{\text{SR}}-G_{\sigma}*S_{\text{bic}},$ (6)
where $G_{\sigma}$ denotes a Gaussian kernel with $\sigma=33$.
A.2 Block-wise Distortion (bd_jup)
The bd_jup feature combines block-wise LPIPS (Zhang et al., 2018) and ERQA (Kirillova et al., 2022) scores. LPIPS is computed on $32\times 32$ blocks with stride 16; ERQA uses $8\times 8$ blocks with no overlap. Since ERQA measures edge-preservation quality (higher is better), we invert it to obtain a distortion score. The final feature is:
$\text{bd\_jup}=0.6\cdot\text{LPIPS}+0.4\cdot(1-\text{ERQA}).$ (7)
Appendix B Crowdsourced annotation dispersion analysis
@@
Appendix C SR fine-tuning scoring and dilation details
@@
Section 5.5 overviews our fine-tuning process to reduce artifacts of existing SR models. Here, we provide more details on the pipeline and on the scoring process.
The input low-resolution image is upscaled with the target SR and RLFN models, and the results are passed to the artifact metric. Then, the regions on the target-SR output where the artifacts were detected are replaced with regions from the RLFN output. The resulting artificial GT image is then used as a target to fine-tune the target-SR model.
@@
  • $\text{Rem}_{img}$ and $\text{Add}_{img}$, image-wise removal and addition rates. The ratio of images where $\left(\left|A\cap B\right|=0\right)\wedge\left(\left|B\right|\neq 0\right)$ determines whether the artifact was removed, and the ratio where $\left|A\cup B\right|>\left|B\right|$ determines whether a new artifact was introduced.
@@
  • $\text{Rem}_{pix}$. Pixel-wise removal rate; represents mean percentage of pixels that were previously classified as artifacts and that go undetected after fine-tuning: $\left|\overline{A}\cap B\right|/\left|B\right|$.
@@
Appendix D Subjective quality of images with artifacts
@@
Appendix E Artifact examples and failure cases
@@
Figure 11 shows examples of prominent artifacts detected by our proposed method across various SR models (Yu et al., 2024; Wang et al., 2021b, 2024a; Liang et al., 2021; Wang et al., 2021a). Each example is annotated with the binary artifact mask and subjective prominence.
Figure 12 shows examples of false detections by our proposed method across SR models (Chen et al., 2023; Hsu et al., 2024; Yue et al., 2023; Wang et al., 2024b; Cai et al., 2019). We observed the following failure cases:
@@
Appendix F Subjective evaluation on additional SR datasets